please help how can I Write the equation of the line passing through each of the given pairs of points. how can I Write the result in slope-intercept form,
(-1, 3) and (4, -2)

Answer by stanbon(48558)

Write the result in slope-intercept form,
(-1, 3) and (4, -2)
----------
slope = [-2-3]/[4--1] = -5/5 = -1
------------------
Slope-intercept form is y = mx+b
You know m=-1, y=3 when x=-1 so you can solve for "b":
3 = (-1)(-1)+b
3 = 1+b
b=2
EQUATION:
y = -x+2
===============
Cheers,
stan H.